# Finding outliers in Gaussian model-based clustering

### v5

@@ 4.3 Crabs Study
Next we evaluate the sensitivity of OCLUST using the crabs dataset (Campbell and Mahon, 1974), which is available in the MASS package (Venables and Ripley, 2002). This study closely mimics the study carried out by Peel and McLachlan (2000) and again by Punzo and McNicholas (2016) to demonstrate their respective approaches to dealing with outliers in model-based clustering. The dataset contains observations for 100 blue crabs, 50 of which are male, and 50 of which are female. The aim for each classification is to recover the sex of the crab. For this study, we will focus on measurements of rear width (RW) and carapace length (CL). We substitute the CL value of the 25th point to one of eight values in $[-15,20]$. The leftmost plot in Figure 5 plots the crabs dataset by sex, with the permuted value in red taking value $\text{CL}=-5$. We use the OCLUST, mean-shift, CNMix, NCM, and DBSCAN algorithms, along with the thresholding approach. For OCLUST, CNMix and NCM, we run each method, restricting the model to one where the clusters have equal shapes and volumes but varying orientations. We let the CNmixt function package determine whether or not the CNMix model was contaminated. Solutions for OCLUST, CNMix, NCM, and the thresholding approach (2T) for the dataset with $\text{CL}=-5$ are plotted in Figure 5. Table 3 summarizes the results for each method, listing the number of misclassified points (M) and the predicted number of outliers ($n_{O}$).
@@
Each method, with the exception of DBSCAN, always identifies the permuted value correctly as an outlier. DBSCAN clusters all the data, including the outlier, into one cluster. NCM classifies more points as outliers as the permuted CL value becomes less extreme. In contrast, OCLUST does the reverse because when the outlier is farther away, it is removed in the gross outlier stage. A gross outlier does not affect the initial clustering, allowing the OCLUST algorithm to remove more points that deviate from multivariate normality. OCLUST identifies more points as outliers than its competitors and has fewer misclassifications in most instances. It is important to note that the outliers found by OCLUST are not simply the misclassifications — i.e., outliers classified as non-outliers and vice versa — of NCM. Instead, as seen in Figure 5, OCLUST identifies three points between the clusters as outliers. This removes the points with high leverage, allowing the clusters to rotate and improve the classification among low values of RW. From one point of view, one may consider that OCLUST removes a few more outliers than necessary; however, their removal improves the parameter estimates. The thresholding approach cannot identify mild outliers because it detects outliers before clustering. Mean shift performs poorly for this dataset due to the shape of the clusters.
@@ 4.4 Wine Data
Finally, we evaluate the results of OCLUST on the wine dataset, available in the gclus R package (Hurley, 2019). This dataset describes 13 attributes of 178 Italian wines (e.g., alcohol, hue, malic acid). Each wine originates from one of three cultivars: Barolo, Barbera, or Grignolino. For this analysis, we add 12 points of noise, uniformly distributed in each dimension as described in Section 4.2. We perform unsupervised classification on this dataset with the aim of removing the noise and recovering the cultivar for each wine. We run OCLUST as well as the five comparator methods from Section 4.3. The results for the comparators, except DBSCAN, are provided in Table 4. DBSCAN performs poorly by classifying all points into one cluster and only identifying seven points of added noise.
@@
NCM achieves good results, with only six misclassifications, five of which are from the Grignolino cultivar. It identifies all 12 points of added noise as outliers. CNMix identifies all 12 outliers and has fewer misclassifications than NCM, all four of which are from the Grignolino cultivar. The thresholding approach performs very well, likely due to the uniform simulation of noise in 13 dimensions creating gross outliers. Mean-shift performs poorly. The solution for OCLUST, with the minimum KL divergence, is shown in Table 4. We have perfect classification among the good points, but 76 points are treated as outliers, which represents 40% of the dataset. Our KL divergence graph in Figure 5(a) has a local minimum at around 35 outliers, which could be viewed as good alternative solution with fewer outliers. Re-running the OCLUST algorithm with our p-value stopping criterion results in it stopping at 35 outliers with p-value 0.059. The classification results with 35 outliers are shown in Table 4. This allows for the preservation of more original data with minimal decrease in classification accuracy. This alternative solution has fewer misclassifications than its comparators. The associated KL divergence plot (Figure 5(a)) displays a pattern unlike Figure 4 because the data are real and not perfectly multivariate normal. We tested each cluster for multivariate normality when all 76 outliers are removed and at the intermediate solution with 35 outliers. We tested the null hypothesis of multivariate normality using the energy test in the energy package (Székely and Rizzo, 2013; Rizzo and Székely, 2022). When 35 outliers are removed, the Barolo, Barbera, and Grignolino clusters have p-values of 0.09, 0.27, and 0.09, respectively. At the minimum KL solution, i.e., with 76 points removed, the Barolo, Barbera, and Grignolino clusters have p-values of 0.77, 0.35, and 0.59, respectively. Evidently, OCLUST continues to remove points until the clusters are unmistakably multivariate normal.
@@
A two-dimensional projection of both OCLUST solutions are shown in Figure 7. In the left-hand figure, where only 35 points are removed, we see some of the identified outliers are far from the clusters but others are mild outliers. In the right-hand plot, we see the solution when more points are removed. Although they may have been correctly classified, they deviate from the symmetric elliptic shape of the multivariate Gaussian distribution. The competing Gaussian model-based classification algorithms, i.e., CNMix and NCM, are able to correctly cluster these points but they seem not to fit well within the model. Identifying these additional outliers may be useful in novelty detection. These wines present inconsistent features which may be relevant to quality control. The spurious points do not fit into the model and indicate wines with chemical and physical properties that differ from the rest.
@@ 5 Discussion
It was shown that, for data from a finite Gaussian mixture model, the log-likelihoods of the subset models are approximately distributed according to a mixture of beta distributions. This result was used to identify outliers, without needing specification of the proportion of outliers, by removing outlying points until the subset log-likelihoods followed this derived distribution. The result is the OCLUST algorithm, which trims outliers from a dataset while clustering using Gaussian mixture models. In simulations, OCLUST performs similarly or better than mean-shift outlier detection for most benchmark datasets except where there is extreme cluster overlap, i.e., S3 and S4. OCLUST performs similarly to NCM when the clusters are elliptical in the real datasets. For the crabs data, OCLUST trims mild outliers with high leverage, which improves the classification for small values of CL. For the wine dataset, OCLUST obtains good classification with fewer iterations. However, if the algorithm is allowed to continue, it removes points which deviate from the assumption of multivariate normality, arguably removing too many points in the process. Although this work used the distribution of the log-likelihoods of the subset models to test for the presence of outliers, the derived distribution may be used to verify other underlying model assumptions, such as whether the clusters are Gaussian. Note that the OCLUST algorithm could be used with other clustering methods and should be effective so long as is it reasonable to assume that the underlying distribution of clusters is Gaussian. Of course, one could extend this work by deriving the distribution of subset log-likelihoods for mixture models with non-Gaussian components. This would allow direct consideration of asymmetric clusters with outliers. Finally, this model is limited by dimensionality because, without sufficient observations, the underlying mixture model becomes over parametrized. One could extend this approach to high-dimensional data by using an analogue of the mixture of factor analyzers model or its extensions (see, e.g., Ghahramani and Hinton, 1997; McNicholas and Murphy, 2008, 2010). Based on the comparisons conducted herein, one might expect the resulting method to perform favourably, or at least comparably, when compared to the approaches used by Wei and Yang (2012) and Punzo et al. (2020).
@@ Acknowledgments
The authors are most grateful to an associate editor and three anonymous reviewers for their helpful comments. This work was supported by an NSERC Undergraduate Research Award, an NSERC Canada Graduate Scholarship, the Canada Research Chairs program, an E.W.R. Steacie Memorial Fellowship, and a Dorothy Killam Fellowship.
@@ Appendix A Relaxing Assumptions
To measure the effect of separation index $J^{*}$ on the approximate log-likelihood $l_{\mathcal{X}}$, 100 random datasets with $n=1800$ for each combination of settings used were generated using the clusterGeneration (Qiu and Joe, 2015) package in R. The settings used were: three clusters with equal cluster proportions, dimensions $p\in\{2,4,6\}$, and separation indices in $[-0.9,0.9]$. Covariance matrices were generated using random eigenvalues $\lambda\in(0,10]$. The parameters were estimated using the EM algorithm. The log-likelihood and the complete-data log-likelihood were calculated using the parameter estimates. The average of the quantity $(l_{\mathcal{X}}-\ell_{\mathcal{X}})/\ell_{\mathcal{X}}$ is computed for each combination of settings — recall that there are 100 datasets for each setting. The result is the average proportional change in log-likelihood over the 100 datasets between the log-likelihood $\ell_{\mathcal{X}}$ and the complete-data log-likelihood $l_{\mathcal{X}}$ (Figure 8).
Figure 8: The effect of cluster separation on the approximate log-likelihood of the model, where the vertical line represents the threshold between separated and overlapping clusters.
@@ B.1 Proof of Lemma 1
$\begin{split}\sum_{g=1}^{G}\pi_{g}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{g},\mathbf{\Sigma}_{g})&=\sum_{g\neq h}\pi_{g}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{g},\mathbf{\Sigma}_{g})+\pi_{h}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{h},\mathbf{\Sigma}_{h})\rightarrow\pi_{h}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{h},\mathbf{\Sigma}_{h}).\end{split}$

### v6

@@ 4.3 Crabs Study
Next we evaluate the sensitivity of OCLUST using the crabs dataset (Campbell and Mahon, 1974), which is available in the MASS package (Venables and Ripley, 2002). This study closely mimics the study carried out by Peel and McLachlan (2000) and again by Punzo and McNicholas (2016) to demonstrate their respective approaches to dealing with outliers in model-based clustering. The dataset contains observations for 100 blue crabs, 50 of which are male, and 50 of which are female. The aim for each classification is to recover the sex of the crab. For this study, we will focus on measurements of rear width (RW) and carapace length (CL). We substitute the CL value of the 25th point to one of eight values in $[-15,20]$. The leftmost plot in Figure 5 plots the crabs dataset by sex, with the permuted value in red taking value $\text{CL}=-5$. We use the OCLUST, mean-shift, CNMix, NCM, and DBSCAN algorithms, along with the thresholding approach. For OCLUST, CNMix and NCM, we run each method, restricting the model to one where the clusters have equal shapes and volumes but varying orientations. We let the CNmixt function determine whether or not the CNMix model is contaminated. Solutions for OCLUST, CNMix, NCM, and the thresholding approach (2T) for the dataset with $\text{CL}=-5$ are plotted in Figure 5. Table 3 summarizes the results for each method, listing the number of misclassified points (M) and the predicted number of outliers ($n_{O}$).
@@
Each method, with the exception of DBSCAN, always identifies the permuted value correctly as an outlier. DBSCAN clusters all the data, including the outlier, into one cluster. NCM classifies more points as outliers as the permuted CL value becomes less extreme. In contrast, OCLUST does the reverse because when the outlier is farther away, it is removed in the gross outlier stage. A gross outlier does not affect the initial clustering, allowing the OCLUST algorithm to remove more points that deviate from multivariate normality.
OCLUST identifies more points as outliers than its competitors and has fewer misclassifications in most instances. It is important to note that the outliers found by OCLUST are not simply the misclassifications — i.e., outliers classified as non-outliers and vice versa — of NCM. Instead, as seen in Figure 5, OCLUST identifies three points between the clusters as outliers. This removes the points with high leverage, allowing the clusters to rotate and improve the classification among low values of RW. From one point of view, one may consider that OCLUST removes a few more outliers than necessary; however, their removal improves the parameter estimates. The thresholding approach cannot identify mild outliers because it detects outliers before clustering. Mean shift performs poorly for this dataset due to the shape of the clusters.
@@ 4.4 Wine Data
Finally, we evaluate the results of OCLUST on the wine dataset, available in the gclus package (Hurley, 2019). This dataset describes 13 attributes of 178 Italian wines (e.g., alcohol, hue, malic acid). Each wine originates from one of three cultivars: Barolo, Barbera, or Grignolino. For this analysis, we add 12 points of noise, uniformly distributed in each dimension as described in Section 4.2. We perform unsupervised classification on this dataset with the aim of removing the noise and recovering the cultivar for each wine. We run OCLUST as well as the five comparator methods from Section 4.3. The results for the comparators, except DBSCAN, are provided in Table 4. DBSCAN performs poorly by classifying all points into one cluster and only identifying seven points of added noise.
@@
NCM achieves good results, with only six misclassifications, five of which are from the Grignolino cultivar. It identifies all 12 points of added noise as outliers. CNMix identifies all 12 outliers and has fewer misclassifications than NCM, all four of which are from the Grignolino cultivar. The thresholding approach performs very well, likely due to the uniform simulation of noise in 13 dimensions creating gross outliers. Mean-shift performs poorly.
The solution for OCLUST, with the minimum KL divergence, is shown in Table 4. We have perfect classification among the good points, but 76 points are treated as outliers, which represents 40% of the dataset. Our KL divergence graph in Figure 6(a) has a local minimum at around 35 outliers, which could be viewed as good alternative solution with fewer outliers. Re-running the OCLUST algorithm with our p-value stopping criterion results in it stopping at 35 outliers with p-value 0.059. The classification results with 35 outliers are shown in Table 4. This allows for the preservation of more original data with minimal decrease in classification accuracy. This alternative solution has fewer misclassifications than its comparators.
The associated KL divergence plot (Figure 6(a)) displays a pattern unlike Figure 4 because the data are real and not perfectly multivariate normal. We tested each cluster for multivariate normality when all 76 outliers are removed and at the intermediate solution with 35 outliers. We tested the null hypothesis of multivariate normality using the energy test in the energy package (Székely and Rizzo, 2013; Rizzo and Székely, 2022). When 35 outliers are removed, the Barolo, Barbera, and Grignolino clusters have p-values of 0.09, 0.27, and 0.09, respectively. At the minimum KL solution, i.e., with 76 points removed, the Barolo, Barbera, and Grignolino clusters have p-values of 0.77, 0.35, and 0.59, respectively. Evidently, OCLUST continues to remove points until the clusters are unmistakably multivariate normal.
@@
Two-dimensional projections of both OCLUST solutions are shown in Figure 7. In the left-hand figure, where only 35 points are removed, we see some of the identified outliers are far from the clusters but others are mild outliers. In the right-hand plot, we see the solution when more points are removed. Although they may have been correctly classified, they deviate from the symmetric elliptic shape of the multivariate Gaussian distribution. The competing Gaussian model-based classification algorithms, i.e., CNMix and NCM, are able to correctly cluster these points but they seem not to fit well within the model. Identifying these additional outliers may be useful in novelty detection. These wines present inconsistent features which may be relevant to quality control. The spurious points do not fit into the model and indicate wines with chemical and physical properties that differ from the rest.
@@ 5 Discussion
It was shown that, for data from a finite Gaussian mixture model, the log-likelihoods of the subset models are approximately distributed according to a mixture of beta distributions. This result was used to identify outliers, without needing specification of the proportion of outliers, by removing outlying points until the subset log-likelihoods followed this derived distribution. The result is the OCLUST algorithm, which trims outliers from a dataset while clustering using Gaussian mixture models. In simulations, OCLUST performs similarly or better than mean-shift outlier detection for most benchmark datasets except where there is extreme cluster overlap, i.e., S3 and S4. OCLUST performs similarly to NCM when the clusters are elliptical in the real datasets. For the crabs data, OCLUST trims mild outliers with high leverage, which improves the classification for small values of CL. For the wine dataset, OCLUST obtains good classification with fewer iterations. However, if the algorithm is allowed to continue, it removes points which deviate from the assumption of multivariate normality, arguably removing too many points in the process.
Although this work used the distribution of the log-likelihoods of the subset models to test for the presence of outliers, the derived distribution may be used to verify other underlying model assumptions, such as whether the clusters are Gaussian. Note that the OCLUST algorithm could be used with other clustering methods and should be effective so long as is it reasonable to assume that the underlying distribution of clusters is Gaussian. Of course, one could extend this work by deriving the distribution of subset log-likelihoods for mixture models with non-Gaussian components. This would allow direct consideration of asymmetric clusters with outliers. Finally, this model is limited by dimensionality because, without sufficient observations, the underlying mixture model becomes over parametrized. One could extend this approach to high-dimensional data by using an analogue of the mixture of factor analyzers model or its extensions (see, e.g., Ghahramani and Hinton, 1997; McNicholas and Murphy, 2008, 2010). Based on the comparisons conducted herein, one might expect the resulting method to perform favourably, or at least comparably, when compared to the approaches used by Wei and Yang (2012) and Punzo et al. (2020).
@@ Acknowledgments
This work was supported by an NSERC Undergraduate Research Award, an NSERC Canada Graduate Scholarship, the Canada Research Chairs program, an E.W.R. Steacie Memorial Fellowship, and a Dorothy Killam Fellowship.
@@ Appendix A Relaxing Assumptions
To measure the effect of separation index $J^{*}$ on the complete-data log-likelihood $l_{\mathcal{X}}$, 100 random datasets with $n=1800$ for each combination of settings used were generated using the clusterGeneration (Qiu and Joe, 2015) package in R. The settings used were: three clusters with equal cluster proportions, dimensions $p\in\{2,4,6\}$, and separation indices in $[-0.9,0.9]$. Covariance matrices were generated using random eigenvalues $\lambda\in(0,10]$. The parameters were estimated using the EM algorithm. The log-likelihood and the complete-data log-likelihood were calculated using the parameter estimates. The average of the quantity $(l_{\mathcal{X}}-\ell_{\mathcal{X}})/\ell_{\mathcal{X}}$ is computed for each combination of settings — recall that there are 100 datasets for each setting. The result is the average proportional change in log-likelihood over the 100 datasets between the log-likelihood $\ell_{\mathcal{X}}$ and the complete-data log-likelihood $l_{\mathcal{X}}$ (Figure 8).
Figure 8: The effect of cluster separation on the complete-data log-likelihood of the model, where the vertical line represents the threshold between separated and overlapping clusters.
@@ B.1 Proof of Lemma 1
$\begin{split}\sum_{g=1}^{G}\pi_{g}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{g},\mathbf{\Sigma}_{g})&=\sum_{g\neq h}\pi_{g}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{g},\mathbf{\Sigma}_{g})+\pi_{h}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{h},\mathbf{\Sigma}_{h})\rightarrow\pi_{h}\phi(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{h},\mathbf{\Sigma}_{h})\end{split}$